### v7

# Decomposed Quadratization: Efficient QUBO Formulation for Learning Bayesian Network

Abstract
Algorithms and hardware for solving quadratic unconstrained binary optimization (QUBO) problems have made significant recent progress. This advancement has focused attention on formulating combinatorial optimization problems as quadratic polynomials. To improve the performance of solving large QUBO problems, it is essential to minimize the number of binary variables used in the objective function. In this paper, we propose a QUBO formulation that offers a bit capacity advantage over conventional quadratization techniques. As a key application, this formulation significantly reduces the number of binary variables required for score-based Bayesian network structure learning. Experimental results on $16$ instances, ranging from $37$ to $223$ variables, demonstrate that our approach requires notably fewer binary variables than quadratization. Moreover, an annealing machine that implement our formulation have outperformed existing algorithms in score maximization.
Figure 1: Overview of our approach to Bayesian network structure learning. We introduce the concept of candidate parent subsets to represent candidate parent sets as a quadratic polynomial.
1 Introduction
A Bayesian network is a probabilistic graphical model that represents a joint probability distribution among random variables in a directed acyclic graph (DAG) (Pearl 1988). One class of associated computational problems is learning the graph structure of a Bayesian network from data. There are two principal approaches to Bayesian network structure learning: constraint-based and score-based (Kitson et al. 2023). Constraint-based algorithms construct graphs through conditional independence tests. Score-based approaches aim to find a DAG with the highest possible score. In this study, we focus on score-based approaches.
Bayesian network structure learning is NP-hard (Chickering, Heckerman, and Meek 2003). Many approaches have been proposed to improve accuracy and reduce execution time. For a small number of variables, some algorithms such as (Cussens 2011) are capable of identifying a DAG with the maximum score. For high-dimensional data, the standard methodology involves the use of approximate approaches. The hill climbing search method over the space of DAGs (Bouckaert 1994) remains competitive despite its simplicity. Approximate searches over the space of topological ordering (Teyssier and Koller 2005) are also well known as competitive algorithms.
Quadratic unconstrained binary optimization (QUBO) is capable of modeling a wide range of combinatorial optimization problems, including image synthesis in the field of computer vision (Kolmogorov and Rother 2007) and various machine learning tasks such as neural networks (Sasdelli and Chin 2021) and decision trees (Yawata et al. 2022). QUBO can also be applied to score-based Bayesian network structure learning (O’Gorman et al. 2014). Note that constraint-based algorithms do not appear to be suitable for QUBO formulation. Recent developments in annealing machines have garnered significant attention for QUBO formulation. An annealing machine is a specialized hardware architecture designed to heuristically solve QUBOs (Yamamoto 2020). In particular, annealing machines based on semiconductor technology derived from simulated annealing have reportedly outperformed existing solvers in maximum cut problems (Matsubara et al. 2020).
The challenge of setting up a QUBO lies in reformulating the objective function into a quadratic polynomial using fewer binary variables. The greater the number of binary variables in a QUBO, the more difficult it generally becomes to solve. More practically, encoding a large number of binary variables onto the hardware circuit of an annealing machine is often infeasible. Quadratization is the common method of QUBO formulation (Heck 2018; Anthony et al. 2016; Boros, Crama, and Heck 2019). Algorithms have been developed to minimize the number of auxiliary variables for quadratization (Verma and Lewis 2020). However, quadratization still requires too many auxiliary variables for certain tasks, including Bayesian network structure learning. Consequently, there is a potential demand for a more efficient QUBO formulation method as an alternative to conventional quadratization.
In this study, we propose a QUBO formulation specifically designed to use fewer binary variables than quadratization. We demonstrate that this formulation offers significant advantages for learning Bayesian networks in terms of bit capacity. Figure 1 illustrates the overview of our approach. Experimental results on $16$ instances, ranging from $37$ to $223$ variables, show that our formulation dramatically reduces the number of binary variables compared to (O’Gorman et al. 2014). These instances can be encoded on the circuit of the Fujitsu Digital Annealer, a fully-coupled annealing machine with $100$K bit capacity (Nakayama et al. 2021). The scores achieved by the Digital Annealer using our formulation were greater than those of existing approaches.
2 Preliminary
We describe a QUBO formulation for learning Bayesian networks, incorporating candidate parent sets into the approach of (O’Gorman et al. 2014). For indexed symbols, we assign the zero index to the empty set and ensure no overlap (e.g., $W_{i0}=\emptyset$ and $\{W_{ij}\}_{j=0}^{\lambda_{i}}$ includes $\lambda_{i}+1$ elements). Appendix A provides the list of symbols.
2.1 Bayesian Network Structure Learning
A Bayesian network, which is a graphical model composed of a DAG and its parameters, represents a joint probability distribution. Each vertex of the DAG corresponds to a random variable. Edges and parameters characterize conditional probability distributions. Let $\mathcal{X}\equiv\{X_{i}\}_{i=1}^{n}$ denote the set of vertices, and $\Pi_{i}$ denote the set of vertices with edges directed towards $X_{i}$. A topological ordering of a graph exists if and only if the graph is a DAG (Cormen et al. 1990).
Definition 2.1.
A topological ordering $\prec$ is a binary relation between any two of vertices such that
  • $X_{i}\prec X_{j}$ and $X_{j}\prec X_{k}\Rightarrow X_{i}\prec X_{k}$.
  • $X_{i}\prec X_{j}\Rightarrow X_{i}\notin\Pi_{j}$.
The goal of score-based Bayesian network structure learning is to find a DAG with the maximum score. A score is the sum of local scores which only rely on the parent set of each vertex. Given complete data, the parent sets $\Pi_{1},\cdots,\Pi_{n}$ are optimized to maximize the score under the constraint that the graph $\mathcal{G}$ corresponding to them is a DAG. This optimization is formulated as
$\mathrm{maximize~{}}\sum_{1\leq i\leq n}\log S_{i}(\Pi_{i})\mathrm{~{}~{}subject~{}to~{}}\mathcal{G}\mathrm{~{}is~{}a~{}DAG},$ (1)
where $S_{i}:2^{\mathcal{X}\setminus\{X_{i}\}}\rightarrow\mathbb{R}$. For simplicity, we take $S_{i}(W)=S_{i}(\emptyset)$ when $|W|$ is greater than the maximum parent set size $m$. The identification of candidate parent sets facilitates narrowing the search space by the relation between parent sets and local scores (de Campos and Ji 2010).
Definition 2.2.
The candidate parent sets of $X_{i}$ are defined as $\{W\subseteq\mathcal{X}\setminus\{X_{i}\}\mid W^{\prime}\subset W\Rightarrow S_{i}(W^{\prime})<S_{i}(W)\}$.
Let $\{W_{ij}\}_{j=0}^{\lambda_{i}}$ denote the candidate parent sets of $X_{i}$, and $\mathcal{X}_{i}$ denote the union of them. Using the candidate parent sets, we can ignore the topological ordering of certain edges that do not produce cycles. Let $\mathcal{E}$ denote the set of edges on possible cycles.
2.2 QUBO Formulation
QUBO is a mathematical optimization problem of minimizing a quadratic polynomial with binary variables. A pseudo-Boolean function maps binary-valued inputs to a real value. Every pseudo-Boolean function can be uniquely represented as a multilinear polynomial given by
$f(\bm{p})\equiv\sum_{V\in\mathcal{F}}\pi(V)\prod_{i\in V}p_{i},$ (2)
where $\bm{p}\equiv(p_{i})_{i=1}^{I}\in\{0,1\}^{I}$, $\pi:2^{\{1,\cdots,I\}}\rightarrow\mathbb{R}\setminus\{0\},\mathcal{F}\subseteq 2^{\{1,\cdots,I\}}\setminus\{\emptyset\}$, and we ignore the constant term. Quadratization is a major technique to convert a higher degree pseudo-Boolean function into a quadratic one using auxiliary variables $\bm{q}\equiv(q_{i})_{i=1}^{J}\in\{0,1\}^{J}$.
Definition 2.3.
The quadratization of $f$ is a quadratic polynomial function $g:\{0,1\}^{I}\times\{0,1\}^{J}\rightarrow\mathbb{R}$ such that
$f(\bm{p})=\min_{\bm{q}\in\{0,1\}^{J}}g(\bm{p},\bm{q}).$ (3)
Rosenberg 1975 has proven that every pseudo-Boolean function can be transformed into a quadratic polynomial through the substitution of the product of two variables with an auxiliary variable and the addition of a penalty term.
2.3 QUBO Formulation for Structure Learning
Bayesian network learning can be uniquely represented by a multilinear polynomial over the state of edges and topological ordering. The state of $\bm{d}\equiv((d_{ij})_{j=1}^{n})_{i=1}^{n}$ is mapped one-to-one to the set of edges, where $d_{ij}=1$ if $X_{j}$ is a parent of $X_{i}$; otherwise, $d_{ij}=0$. The score component is
$O(\bm{d})\equiv\sum_{1\leq i\leq n}\sum_{W\subseteq\mathcal{X}_{i}}\pi_{i}(W)\prod_{1\leq j\leq n;X_{j}\in W}d_{ij},$ (4)
where $\pi_{i}(W)\equiv\sum_{W^{\prime}\subseteq W}(-1)^{1+|W\setminus W^{\prime}|}\log S_{i}(W^{\prime})$. The higher-degree terms in the score component can be transformed into quadratic terms through quadratization. The state of $\bm{r}\equiv(r_{ij})_{(i,j)\in\mathcal{E}}$ corresponds to the topological ordering, where $r_{ij}=0$ if the order of $X_{j}$ is higher than $X_{i}$; otherwise, $r_{ij}=1$. The linear ordering of vertices and the consistency of edges in definition 2.1 are captured by
$C(\bm{d},\bm{r})\equiv$ $\sum_{(i,j),(j,k),(i,k)\in\mathcal{E}}\delta_{1}R(r_{ij},r_{jk},r_{ik})$
$+\sum_{(i,j)\in\mathcal{E}}\delta_{2}\bigl{(}d_{ij}r_{ij}+d_{ji}(1-r_{ij})\bigr{)},$ (5)
where $\delta_{1},\delta_{2}\in(0,\infty)$ and $R(r_{ij},r_{jk},r_{ik})\equiv r_{ik}+r_{ij}r_{jk}-r_{ij}r_{ik}-r_{jk}r_{ik}$. There exist $(\delta_{1},\delta_{2})$ such that $C(\bm{d},\bm{r})=0$ when the objective function $O(\bm{d})+C(\bm{d},\bm{r})$ is minimized. Additionally, the graph corresponding to the state of $\bm{d}$ is a DAG if $C(\bm{d},\bm{r})=0$. Consequently, this QUBO formulation is equivalent to learning Bayesian networks.
3 Decomposed Quadratization
Searching over candidate parent sets is a key technique in Bayesian network structure learning. To deal with candidate parent sets in a QUBO formulation, we use auxiliary variables corresponding to the sets of primary variables. The auxiliary variables capture the score component and identify the state of primary variables. We decompose a quadratization into the optimization of auxiliary variables and the transformation that maps them to primary variables. This decompostion of quadratization is defined as
Definition 3.1.
The decomposed quadratization of $f$ is the pair of a function $g_{1}:\{0,1\}^{J}\rightarrow\{0,1\}^{I}$ and a quadratic polynomial function $g_{2}:\{0,1\}^{J}\rightarrow\mathbb{R}$ such that
$f(\bm{p})=\min_{\bm{q}\in\{0,1\}^{J};\bm{p}=g_{1}(\bm{q})}g_{2}(\bm{q}).$ (6)
Example 3.2.
Let $g_{1}(q_{1},q_{2})=(q_{1},q_{2},q_{2})$ and $g_{2}(q_{1},q_{2})=-q_{1}-3q_{2}+2q_{1}q_{2}$. Then $(g_{1},g_{2})$ is a decomposed quadratization of $f(p_{1},p_{2},p_{3})=-p_{1}-3p_{2}p_{3}+2p_{1}p_{2}p_{3}$.
Here we describe the behavior of decomposed quadratization. Considering the case where each $p_{i}$ corresponds to $q_{i}$, the concept of decomposed quadratization encompasses quadratization. To capture the subsets of the terms in $\mathcal{F}$, we make a set $\emptyset\subset Q_{i}\subseteq\{1,\cdots,I\}$ correspond to a binary variable $q_{i}\in\{0,1\}$ one-to-one. Let $\bar{g}_{1},\bar{g}_{2}$ denote
$\bar{g}_{1}(\bm{q})\equiv(\min\{1,\sum_{1\leq j\leq J;i\in Q_{j}}q_{j}\})_{i=1}^{I},$ (7)
$\bar{g}_{2}(\bm{q})\equiv\sum_{1\leq i\leq j\leq J}\bar{\pi}(Q_{i},Q_{j})q_{i}q_{j}+h(\bm{q}),$ (8)
where $\bar{\pi}:2^{\{1,\cdots,I\}}\times 2^{\{1,\cdots,I\}}\rightarrow\mathbb{R}$, and $h:\{0,1\}^{J}\rightarrow[0,\infty)$is a quadratic polynomial. To discuss the behavior of $(\bar{g}_{1},\bar{g}_{2})$, we introduce theorem 3.5. Assumption 3.3 ensures the comprehensiveness of the terms in $\mathcal{F}$. Assumption 3.4 presents the inclusion relations among $\{Q_{i}\}_{i=1}^{J}$.
Assumption 3.3.
$\mathcal{F}\subseteq\bigcup_{1\leq i\leq j\leq J}\{Q_{i}\cup Q_{j}\}$.
Assumption 3.4.
If there exists a pair of an index $1\leq i\leq J$ and a set $\mathcal{W}\subseteq\{1,\cdots,J\}\setminus\{i\}$ with $|\mathcal{W}|\geq 2$ such that
$Q_{i}\subseteq\bigcup_{j\in\mathcal{W}}Q_{j}\mathrm{~{}and~{}}Q_{i}\nsubseteq\bigcup_{j\in\mathcal{W}^{\prime}}Q_{j}\mathrm{~{}for~{}any~{}}\mathcal{W}^{\prime}\subset\mathcal{W},$ (9)
then $Q_{i}=Q_{j}\cup Q_{k}$ for some $j,k\in\{1,\cdots,J\}\setminus\{i\}$.
Theorem 3.5.
Suppose that assumption 3.3 and assumption 3.4 hold. Then there exists $(\bar{\pi},h)$ such that $(\bar{g}_{1},\bar{g}_{2})$ is a decomposed quadratization of $f$.
Proof.
See section B.1. ∎
To capture the inclusion relations, we consider designing a penalty function $h$ that induces
$Q_{i}\subseteq\bigcup_{j\in\mathcal{W}}Q_{j}\Rightarrow(1-q_{i})\prod_{j\in\mathcal{W}}q_{j}=0.$ (10)
According to lemma 3.6, one element cannot be directly covered by more than three others. This prevents the direct incorporation of eq. 10 into a penalty. Assumption 3.4 indirectly addresses it by covering one element with two others. Lemma 3.7 restricts the formulation of penalties to capture this assumption.
Lemma 3.6.
For $K\in\{3,4,\cdots\}$, no quadratic polynomial function $\phi:\{0,1\}^{K+1}\rightarrow\mathbb{R}$ satisfies
$\phi(q_{1},\cdots,q_{K},0)=0\Leftrightarrow\prod_{1\leq i\leq K}q_{i}=0.$ (11)
Proof.
See section B.2. ∎
Lemma 3.7.
Suppose that a quadratic polynomial function $\phi:\{0,1\}^{3}\rightarrow\mathbb{R}$ satisfies eq. 11. Then the following holds:
$\phi(0,0,1)+\phi(1,1,1)\neq\phi(1,0,1)+\phi(0,1,1).$ (12)
Proof.
See section B.3. ∎
Quadratization is a specific form of decomposed quadratization that requires assumption 3.8 in addition to assumption 3.3 and assumption 3.4. This additional assumption of quadratization is disadvantageous for reducing the number of auxiliary variables in certain tasks such as example 3.9.
Assumption 3.8.
$\{i\}\in\{Q_{j}\}_{j=1}^{J}$ for all $1\leq i\leq I$.
Example 3.9.
Let =$\mathcal{F}=\{\{1,2,3,4\},\{1,2,3\},\{4\},\{3,4\}\}$. A decomposed quadratization requires $3$ auxiliary variables: $Q_{1}=\{1,2,3\},Q_{2}=\{4\}$, and $Q_{3}=\{3\}$. A quadratization requires $6$ auxiliary variables: $Q_{1}=\{1\},Q_{2}=\{2\},Q_{3}=\{3\},Q_{4}=\{4\},Q_{5}=\{1,2\}$, and $Q_{6}=\{3,4\}$.
Considering the biases and variable couplers in a QUBO, the auxiliary variables satisfy $J\geq\lceil-\frac{1}{2}+\frac{1}{2}\sqrt{1+8|\mathcal{F}|}\rceil$. We minimize the number of auxiliary variables to capture assumption 3.3 and assumption 3.4. An optimal solution of $\{Q_{i}\}_{i=1}^{J}$ is a subset of $\bigcup_{V\in\mathcal{F}}2^{V}\setminus\{\emptyset\}$. Let $V_{\mathrm{close}},V_{\mathrm{open}}\subseteq\bigcup_{V\in\mathcal{F}}2^{V}\setminus\{\emptyset\}$ such that $V_{\mathrm{close}}\cap V_{\mathrm{open}}=\emptyset$, $V_{\mathrm{close}}$ is a subset of an optimal solution, and $V_{\mathrm{close}}\cup V_{\mathrm{open}}$ contains an optimal solution. Since $V_{\mathrm{close}}$ is a subset of an optimal solution, we can ignore the terms in $\mathcal{F}$ that can be represent by the union of some two elements in $V_{\mathrm{close}}$. Let $W_{\mathrm{open}}\subseteq\mathcal{F}$ be the terms that still cannot be ignored. We provide algorithm 1 to find an optimal solution. The mechanism of this algorithm is as follows:
Row 2:
If there is only one way to represent $W$, then $V$ is included in an optimal solution.
Row 3:
We can ignore elements in $W_{\mathrm{open}}$ represented by pairs of elements in $V_{\mathrm{close}}$.
Row 4:
If $V$ is not a subset of any element in $W_{\mathrm{open}}$, then it is not included in an optimal solution.
Here we minimize $\sum_{i=1}^{\nu}v_{i}$ under the constraint that assumption 3.3 and assumption 3.4 hold for
$\{Q_{i}\}_{i=1}^{J}=V_{\mathrm{close}}\cup\{V_{i}\mid v_{i}=1\}_{i=1}^{\nu},$ (13)
where $\bm{v}\equiv(v_{i})_{i=1}^{\nu}\in\{0,1\}^{\nu}$ and $\{V_{i}\}_{i=1}^{\nu}\equiv V_{\mathrm{open}}$. By expressing the constraints as linear inequalities, the search for an optimal solution can be formulated as an integer linear programming problem of $\bm{v}$. See section C.1. This formulation, with some conversion, can be adapted to minimize the number of terms. See section C.2.
Algorithm 1 Search Space Reduction
1:  $W_{\mathrm{open}}\leftarrow\mathcal{F},V_{\mathrm{close}}\leftarrow\emptyset,V_{\mathrm{open}}\leftarrow\bigcup_{V\in\mathcal{F}}2^{V}\setminus\{\emptyset\}$.
2:  Merge $\{V\in V_{\mathrm{open}}\mid$ for some $W\in W_{\mathrm{open}}$, $V$ is only an element in $V_{\mathrm{open}}$ such that $W\in\bigcup_{V^{\prime}\in\{\emptyset\}\cup V_{\mathrm{close}}}\{V\cup V^{\prime}\}\}$ into $V_{\mathrm{close}}$.
3:  Exclude $\bigcup_{V,V^{\prime}\in V_{\mathrm{close}}}\{V\cup V^{\prime}\}$ from $W_{\mathrm{open}}$.
4:  Exclude $V_{\mathrm{close}}\cup\{V\in V_{\mathrm{open}}\mid$ there does not exist $W\in W_{\mathrm{open}}$ such that $V\subseteq W\}$ from $V_{\mathrm{open}}$.
Figure 2: Example of the QUBO formulation with candidate parent subsets. An undirected edge between two binary variables represents that the objective function contains their product term. Gray signifies that the state of a binary variable is 00 and white is $1$. Each color corresponds to the edge in the DAG. Let $n=3,m=2,\lambda_{1}=3,\lambda_{2}=1,\lambda_{3}=1,\Pi_{1}=\{X_{2}\},\Pi_{2}=\emptyset,\Pi_{3}=\{X_{1},X_{2}\},W_{11}=\{X_{2}\},W_{12}=\{X_{3}\},W_{13}=\{X_{2},X_{3}\},W_{21}=\{X_{1},X_{3}\},W_{31}=\{X_{1},X_{2}\},\mu_{1}=2,\mu_{2}=1,\mu_{3}=1,U_{11}=\{X_{2}\},U_{12}=\{X_{3}\},U_{21}=\{X_{1},X_{3}\}$, and $U_{31}=\{X_{1},X_{2}\}$.
4 Efficient QUBO for Structure Learning
In this section, we propose the decomposed quadratization of $O$ to encode candidate parent sets efficiently. Here we introduce the concept of candidate parent subsets.
Definition 4.1.
The candidate parent subsets of $X_{i}$ are defined as the set containing elements $U,U^{\prime}\subseteq\mathcal{X}_{i}$ such that $U\cup U^{\prime}=W$, for any $W$ in the candidate parent sets of $X_{i}$.
Let $\{U_{ij}\}_{j=0}^{\mu_{i}}$ denote the candidate parent subsets of $X_{i}$. To encode candidate parent subsets in a QUBO, we make a candidate parent subset $U_{ij}$ correspond to a binary variable $u_{ij}\in\{0,1\}$ one-to-one. The state of $\bm{d}$ is identified by
$D(\bm{u})\equiv((\min\{1,\sum_{1\leq k\leq\mu_{i};X_{j}\in U_{ik}}u_{ik}\})_{j=1}^{n})_{i=1}^{n},$ (14)
where $\bm{u}\equiv((u_{ij})_{j=1}^{\mu_{i}})_{i=1}^{n}$. A parent set is captured by the union of at most two elements included in candidate parent subsets. Since each vertex has just one parent set, we can disregard the inclusion relations among candidate parent subsets as mentioned in assumption 3.4. From definition 4.1, the following proposition equivalent to assumption 3.3 holds.
Proposition 4.2.
$\mathcal{F}_{i}\equiv\{W_{ij}\}_{j=1}^{\lambda_{i}}\subseteq\bigcup_{1\leq j\leq k\leq\mu_{i}}\{U_{ij}\cup U_{ik}\}\subseteq 2^{\mathcal{X}_{i}}\setminus\{\emptyset\}$ for all $1\leq i\leq n$.
The main distinction of our approach from section 2.3 is that the objective function is represented not by edges among vertices but by candidate parent subsets. To enforce exactly one parent set per vertex (i.e., $\sum_{1\leq j\leq\mu_{i}}u_{ij}\leq 2$), we use the constraint characterized by a penalty coefficient $\xi\in(0,\infty)$ and auxiliary variables $\bm{z}=(z_{i})_{i=1}^{n}\in\{0,1\}^{n}$. Then the score component is
$\bar{O}(\bm{u},\bm{z})\equiv\sum_{1\leq i\leq n}\sum_{1\leq j\leq k\leq\mu_{i}}\bar{\pi}_{i}(U_{ij},U_{ik})u_{ij}u_{ik}$
$+\sum_{1\leq i\leq n}\xi\Bigl{(}z_{i}-z_{i}\sum_{1\leq j\leq\mu_{i}}u_{ij}+\sum_{1\leq j<k\leq\mu_{i}}u_{ij}u_{ik}\Bigr{)},$ (15)
where $\bar{\pi}_{i}(U,U^{\prime})\equiv(-1)^{1+|\{U\}\cup\{U^{\prime}\}|}(\log S_{i}(U\cup U^{\prime})-\log S_{i}(U)-\log S_{i}(U^{\prime})+\log S_{i}(\emptyset))$. Using proposition 4.2, our formulation requires equal to or fewer binary variables than the quadratization of $O$. For candidate parent subsets in quadratization, it is essential to capture the relation between them and the edges as follows:
$(1-u_{ij})\prod_{1\leq k\leq n;X_{k}\in U_{ij}}d_{ik}=0.$ (16)
Lemma 3.6 demonstrates that such relation cannot be captured by a quadratic polynomial when a candidate parent subset includes more than three elements. Consequently, we benefit from the use of decomposed quadratization with candidate parent subsets. The following lemma provides a lower bound for the penalty coefficient $\xi$.
Lemma 4.3.
Let $\xi_{0}$ be defined as
$\xi_{0}\equiv-3\min_{1\leq i\leq n}\min_{0\leq j\leq k\leq\mu_{i}}\bar{\pi}_{i}(U_{ij},U_{ik}).$ (17)
Suppose that $\xi>\xi_{0}$ holds and the state of $\bm{d}$ aligns with candidate parent sets. Then $(D,\bar{O})$ is a decomposed quadratization of $O$.
Proof.
See section B.4. ∎
The constraint for definition 2.1 is also captured by candidate parent subsets. Let $\bar{C}(\bm{u},\bm{r})$ denote the right side of section 2.3 where $d_{ij}$ is replaced by
$\sum_{1\leq k\leq\mu_{i};X_{j}\in U_{ik}}u_{ik},$ (18)
where this substitution does not use additional auxiliary variables, but may increase the number of terms. The objective function in our approach is $\bar{H}(\bm{u},\bm{r},\bm{z})\equiv\bar{O}(\bm{u},\bm{z})+\bar{C}(\bm{u},\bm{r})$. Figure 2 illustrates this formulation. The following theorem guarantees that optimal networks can be obtained using our QUBO formulation.
Theorem 4.4.
Let $\delta_{0}$ be defined as
$\delta_{0}\equiv\max_{1\leq i\leq n,0\leq j\leq\lambda_{i}}(\log S_{i}(W_{ij})-\log S_{i}(\emptyset)).$ (19)
Suppose that $\xi>\xi_{0}$ and $\delta_{2}>(n-2)\delta_{1}>(n-2)\delta_{0}$ hold. Then $\bar{C}(\bm{u},\bm{r})=0$ holds when $\bar{H}(\bm{u},\bm{r},\bm{z})$ is minimized.
Proof.
See section B.5. ∎
Algorithm 2 Candidate Parent Subset Reduction
1:  $W_{\mathrm{open}}\leftarrow\mathcal{F}_{i},V_{\mathrm{close}}\leftarrow\emptyset,V_{\mathrm{open}}\leftarrow\bigcup_{V\in\mathcal{F}_{i}}2^{V}\setminus\{\emptyset\}$.
2:  while $V_{\mathrm{close}}$ has changed since the last evaluation, do
3:     Exclude $\{V\in V_{\mathrm{open}}\mid V\subseteq W$ for at most one element $W\in W_{\mathrm{open}}\}$ from $V_{\mathrm{open}}$.
4:     Execute the row from $2$ to $4$ in algorithm 1.
5:     Merge $W_{\mathrm{open}}\setminus\bigcup_{V,V^{\prime}\in V_{\mathrm{close}}\cup V_{\mathrm{open}}}\{V\cup V^{\prime}\}$ into $V_{\mathrm{close}}$, and exclude it from $W_{\mathrm{open}}$.
6:  end while
7:  Merge $W_{\mathrm{open}}\setminus\bigcup_{V\in V_{\mathrm{open}},V^{\prime}\in V_{\mathrm{close}}}\{V\cup V^{\prime}\}$ into $V_{\mathrm{open}}$.
The number of required binary variables is $\sum_{1\leq i\leq n}(\mu_{i}+1)+|\mathcal{E}|$. We find candidate parent subsets to minimize $\mu_{i}$ under the constraint to capture proposition 4.2. An optimal solution of candidate parent subsets of $X_{i}$ is a subset of $\bigcup_{V\in\mathcal{F}_{i}}2^{V}\setminus\{\emptyset\}$. Since assumption 3.4 is not required, algorithm 2 can reduce the search space of candidate parent subsets. This algorithm works as follows:
Row 3:
We temporarily ignore elements that contribute to represent at most one element included in $W_{\mathrm{open}}$.
Row 5:
Considering the removed elements at row 3, an optimal solution includes the elements in $W_{\mathrm{open}}$ that cannot be represented by pairs of elements in $V_{\mathrm{close}}\cup V_{\mathrm{open}}$.
Row 7:
From a perspective similar to row 5, elements in $W_{\mathrm{open}}$ that are represented solely by pairs of elements in $V_{\mathrm{open}}$ may be included in an optimal solution.
Additionally, we can exclude the constraint that captures assumption 3.4 from the integer linear programming problem.
[] Instance $m$ O’Gorman et al. IP only IP + algorithm 2 Number Time [s] Number Time [s] Number Time [s] alarm $4$ $3982\pm 346$ $272\pm 61$ $\bm{1373\pm 46}$ $61\pm 29$ $\bm{1373\pm 46}$ $\bm{14\pm 2}$ barley $4$ $1826\pm 279$ $271\pm 149$ $\bm{995\pm 143}$ $\bm{11\pm 2}$ $\bm{995\pm 143}$ $\bm{11\pm 2}$ hailfinder $4$ $2374\pm 31$ $52\pm 24$ $\bm{1913\pm 6}$ $\bm{4\pm 1}$ $\bm{1913\pm 6}$ $\bm{4\pm 1}$ hepar2 $4$ $3629\pm 423$ $51\pm 16$ $\bm{2616\pm 133}$ $3\pm 0$ $\bm{2616\pm 133}$ $\bm{2\pm 0}$ chess $2$ $7040\pm 167$ $\bm{63\pm 2}$ $\bm{6964\pm 165}$ $89\pm 4$ $\bm{6964\pm 165}$ $103\pm 4$ chess $3$ $80116\pm 3014$ $\bm{3229\pm 151}$ $26355\pm 1403$ $7268\pm 492$ $\bm{23662\pm 1103}$ $5567\pm 376$ win95pts $2$ $5150\pm 39$ $\bm{27\pm 2}$ $\bm{5028\pm 42}$ $33\pm 3$ $\bm{5028\pm 42}$ $37\pm 4$ win95pts $3$ $37705\pm 1935$ $1223\pm 90$ $9220\pm 398$ $1036\pm 117$ $\bm{9025\pm 333}$ $\bm{890\pm 105}$ pathfinder $2$ $10831\pm 382$ $\bm{178\pm 8}$ $\bm{10725\pm 379}$ $259\pm 30$ $\bm{10725\pm 379}$ $284\pm 20$ pathfinder $3$ $107725\pm 5329$ $\bm{11695\pm 834}$ $38245\pm 3242$ $30707\pm 4322$ $\bm{34448\pm 3016}$ $21364\pm 3773$ mushroom $2$ $18050\pm 508$ $\bm{318\pm 13}$ $\bm{17930\pm 506}$ $681\pm 32$ $\bm{17930\pm 506}$ $513\pm 23$ connect $2$ $15155\pm 334$ $\bm{137\pm 9}$ $\bm{14978\pm 340}$ $228\pm 17$ $\bm{14978\pm 340}$ $240\pm 12$ connect $3$ $179788\pm 6242$ $\bm{9897\pm 602}$ $64533\pm 4352$ $30381\pm 4267$ $\bm{58247\pm 4053}$ $20106\pm 2661$ munin1 $2$ $25180\pm 266$ $\bm{699\pm 20}$ $\bm{25001\pm 268}$ $953\pm 75$ $\bm{25001\pm 268}$ $1079\pm 67$ andes $2$ $29502\pm 171$ $\bm{40\pm 1}$ $\bm{28810\pm 182}$ $249\pm 37$ $\bm{28810\pm 182}$ $250\pm 35$ andes $3$ $92838\pm 3735$ $2182\pm 289$ $\bm{31162\pm 136}$ $504\pm 132$ $\bm{31162\pm 136}$ $\bm{217\pm 8}$
Table 1: Number of binary variables and execution time for the QUBO formulation. The mean and standard deviation of $5$ trials are presented. All the candidate parent subsets not proven optimal within $60$ [s] were feasible solutions. The bold values represent the smallest number of binary variables and the shortest execution times among all approaches.
5 Experimental Results
In this section, we investigate the performance of our approach for reducing the required number of binary variables. We also demonstrate the score maximization using some solvers with our QUBO formulation. Version 9.1.2 of the Gurobi Optimizer  [^1] https://www.gurobi.com/ was used to solve the integer linear programming problems. The fourth-generation Fujitsu Digital Annealer was employed for score maximization. All experiments, except for those using the Digital Annealer, were conducted on a 64-bit Windows machine with an Intel Xeon W-2265 @ $3.50$ GHz and $128$ GB of RAM. The code was implemented in the Julia programming language 1.5.3 version. As benchmarks, we adopted $5$ simulated datasets from each of the $8$ discrete networks: alarm ($n=37$), barley ($n=48$), hailfinder ($n=56$), hepar2 ($n=70$), win95pts ($n=76$), pathfinder ($n=109$), munin1 ($n=186$), and andes ($n=223$) from the Bayesian network repository  [^2] https://www.bnlearn.com/bnrepository/. In addition, we used $5$ times non-recoverable extracted datasets from each of the $3$ datasets: chess ($n=75$), mushroom ($n=117$), and connect ($n=129$) from the Frequent Item Set Mining Dataset Repository  [^3] http://fimi.uantwerpen.be/data/. The sample size of each dataset is $1000$.
5.1 QUBO Formulation
We show that our formulation reduces the number of binary variables required in a QUBO.
Settings.  The objective function was $\bar{H}(\bm{u},\bm{r},\bm{z})$. The score function was the BDeu score (Buntine 1991), where the equivalent sample size was set to $1$. See section D.1. The identification of candidate parent sets was performed exactly. Candidate parent sets were identified by simply enumerating all the elements in $\{W\subseteq\mathcal{X}\setminus\{X_{i}\}\mid|W|\leq m\}$ and comparing their local scores when the inclusion relations between them was in place. The maximum parent set size $m$ for each instance with more than $75$ variables was set to the largest value that allowed the identification process to be completed within $48$ [h]. The number of candidate parent sets and the execution time for their identification are found in section D.2. The integer linear programming problem to find optimal candidate parent subsets was formulated as section C.1. The time limit for each variable was $60$ [s]. The execution time for the QUBO formulation includes that of algorithm 2, setting up the integer linear programming problem, solving it, and incorporating the solution into the objective function.
Baselines.  We compare our formulation with the approach proposed by (O’Gorman et al. 2014). Each vertex requires $\lfloor\frac{1}{4}(|\mathcal{X}_{i}|-1)^{2}\rfloor+2$ auxiliary variables when $m_{i}=3$, and $\frac{1}{2}|\mathcal{X}_{i}|(|\mathcal{X}_{i}|-1)+3$ when $m_{i}=4$, where $m_{i}\equiv\max_{0\leq j\leq\lambda_{i}}|W_{ij}|$. To investigate the merit of algorithm 2, we found candidate parent subsets using the integer linear programming problem with $W_{\mathrm{open}}=\mathcal{F}_{i},V_{\mathrm{open}}=\bigcup_{V\in\mathcal{F}_{i}}2^{V}\setminus\{\emptyset\}$, and $V_{\mathrm{close}}=\emptyset$. We refer to this as ”IP only” and our approach with algorithm 2 as ”IP + algorithm 2”. The candidate parent sets to formulate these baselines were the same as those in the previously described setting. The execution time for each baseline does not include the time spent on identifying candidate parent sets.
Evaluation.  The comparison between IP + algorithm 2 and IP only in table 1 demonstrates that optimizing candidate parent subsets with algorithm 2 is advantageous in terms of both the number of binary variables and execution time when the parent set size is larger. The quadratization for some instances with $m=3$ required binary variables over the $100$K bit capacity of the Digital Annealer. The decomposed quadratization in our approach significantly reduced the required number of binary variables compared to the quadratization approach. The execution time of our formulation was disadvantageous for most larger networks, primarily due to the increasing number of terms. Table 2 shows that incorporating the candidate parent subsets into the QUBO introduced a significant bottleneck.
5.2 Score Maximization
We demonstrate score maximization using the QUBO formulation by IP + algorithm 2.
Settings.  The time limit for each trial was set to $3600$ [s]. The coefficients in the penalty terms were $\delta_{1}=1.1\delta_{0},\delta_{2}=1.1(n-2)\delta_{1}$, and $\xi=1.1\xi_{0}$. We used two heuristic solvers and one exact method: the Digital Annealer (DAQ), classical simulated annealing (SAQ), and the Gurobi Optimizer (GOQ). The inequality constraint is placed outside the objective function in both DAQ and GOQ. This formulation is strictly categorized as binary quadratic programming. See section D.4. The parameters for DAQ and GOQ were set to their default values. The annealing schedule of SAQ was geometric, starting with $T_{0}=100$. See section D.3. Additional bits for a minor embedding (Choi 2008, 2010; Eppstein 2009) were not required because the Digital Annealer is a fully-coupled type. Note that the short description of minor embedding can be found in section D.6.
Baselines.  The baseline algorithms are only score-based. We adopted three approximate approaches and one exact algorithm: the hill climbing search method (HCS), the simulated annealing over ordering space (SAO), the acyclic selection ordering-based search (ASO), and the GOBNILP software  [^4] https://www.cs.york.ac.uk/aig/sw/gobnilp/ (GOB). The HCS and SAO are approximate approaches described in (Scutari, Graafland, and Gutiérrez 2019). The ASO and GOB are known as competitive algorithms. The tabu list length in HCS was set to $10$, and the state of $10$ randomly selected possible edges was changed upon termination of the greedy search. The annealing schedule of SAO was same as that of SAQ. The ASO was the version proposed in (Scanagatta et al. 2015). The version of GOB was the pygobnilp1.0 that relies on the Gurobi Optimizer. We used common candidate parent sets in our formulation, except for HCS. To ensure a fair comparison, the time limit for each of SAO, ASO, and GOB was set at $3600$ [s] plus the execution time required for the QUBO formulation by IP + algorithm 2. The time limit for HCS was further extended by the time needed to identify candidate parent sets. We ignored the time to transfer the encoding information online to the Digital Annealer environment. The HCS and ASO were restarted repeatedly during the time limit.
Evaluation.  Table 3 displays the results of score maximization. For the smaller instances, GOB identified optimal networks. In most other cases, the scores achieved by DAQ exceeded those of the baselines. If the long execution time required for the QUBO formulation can be reduced in future work, the difference will increase further. The Digital Annealer effectively highlighted the strengths of our formulation compared to the results from SAQ and GOQ using the same QUBO formulation.
[] Instance $m$ Time [s] alarm $4$ $1\pm 0,3\pm 1,10\pm 1$ barley $4$ $0\pm 0,0\pm 0,11\pm 2$ hailfinder $4$ $0\pm 0,0\pm 0,4\pm 1$ hepar2 $4$ $0\pm 0,0\pm 0,2\pm 0$ chess $2$ $2\pm 0,0\pm 0,100\pm 4$ chess $3$ $406\pm 49,2554\pm 199,2608\pm 195$ win95pts $2$ $1\pm 0,0\pm 0,36\pm 4$ win95pts $3$ $45\pm 5,398\pm 82,446\pm 54$ pathfinder $2$ $3\pm 0,0\pm 0,281\pm 19$ pathfinder $3$ $1407\pm 272,1679\pm 184,18278\pm 3349$ mushroom $2$ $23\pm 1,0\pm 0,490\pm 22$ connect $2$ $5\pm 0,0\pm 0,235\pm 12$ connect $3$ $2472\pm 436,4129\pm 148,13505\pm 2114$ munin1 $2$ $12\pm 1,0\pm 0,1067\pm 67$ andes $2$ $1\pm 0,0\pm 0,249\pm 35$ andes $3$ $6\pm 1,2\pm 1,209\pm 8$
Table 2: Execution time for IP + algorithm 2. The execution time for each of the following three steps is displayed. Left : algorithm 2 and preparation of the interger linear programming problem. Middle : finding candidate parent subsets by solving the problem. Right : incorporating the candidate parent subsets into the objective function of the QUBO. Note that the execution time for IP only is shown in section D.5.
[] Instance $m$ DAQ GOQ HCS SAO GOB alarm $4$ 00 00 00 00 $\bm{5}$ barley $4$ 00 00 00 00 $\bm{5}$ hailfinder $4$ 00 00 00 00 $\bm{5}$ hepar2 $4$ 00 00 00 00 $\bm{5}$ chess $2$ $\bm{5}$ 00 00 00 00 chess $3$ $\bm{5}$ 00 00 00 00 win95pts $2$ $1$ 00 00 00 $\bm{4}$ win95pts $3$ $2$ 00 00 $\bm{3}$ 00 pathfinder $2$ $\bm{4}$ 00 00 $1$ 00 pathfinder $3$ $\bm{5}$ 00 00 00 00 mushroom $2$ $\bm{5}$ 00 00 00 00 connect $2$ $\bm{5}$ 00 00 00 00 connect $3$ $\bm{2}$ $1$ $1$ $1$ 00 munin1 $2$ $\bm{5}$ 00 00 00 00 andes $2$ $\bm{5}$ 00 00 00 00 andes $3$ $\bm{5}$ 00 00 00 00
Table 3: Results of score maximization. The number of trials achieving the highest score is displayed. The highest score refers to the shortest execution time when the scores are the same. For the four instances above the line, the GOB identified optimal networks within a few seconds. While the GOQ identified an optimal network in only one trial of the barley instance, the execution time was longer than the GOB. The solutions for the win95pts with $m=2$ and the connect with $m=3$ were not proven to be optimal. The SAQ and ASO never won in any of the trials. The bold values show the greatest number of wins among all solvers.
6 Conclusion
We proposed a QUBO formulation tailored to score-based Bayesian network structure learning. The essence of this approach lies in reducing the number of required binary variables through decomposed quadratization with candidate parent subsets. We also provided an algorithm to efficiently find optimal candidate parent subsets. Experimental results demonstrated that our approach significantly reduced the number of binary variables compared to the previous work based on quadratization. Additionally, our formulation using the Digital Annealer achieved improved BDeu scores over existing methods for medium-sized instances. We expect that our approach can be more effectively applied to larger-scale structure learning problems in the future development of annealing processors.
Limitations.  We have identified three limitations in our approach. First, the execution time to incorporate the candidate parent subsets into the objective function is a potential drawback for larger-scale problems. Secondly, capturing a parent set with two bits may be disadvantageous for a 1-bit inversion search, as transitioning from one candidate parent set to another necessitates an inversion of up to four bits. Lastly, our formulation could potentially have a negative impact by increasing the number of terms in a QUBO. In particular, a minor embedding requires additional bits as the number of terms increases.
Future Works.  Along with further investigation addressing the above limitations, we will explore the application of decomposed quadratization to tasks beyond Bayesian network structure learning. Decomposed quadratization can serve as a promising alternative to conventional quadratization for a wide range of tasks involving multilinear polynomials with higher-degree terms.
Acknowledgments
We acknowledge Fujitsu Limited for their support in providing access to the fourth-generation Fujitsu Digital Annealer.
Appendix A Notations
[] Notation Description $n$ the number of vertices $m$ the maximum parent set size $\mathcal{X}$ $\mathcal{X}\equiv\{X_{i}\}_{i=1}^{n}$ is the set of vertices $\bm{d}$ $\bm{d}\equiv((d_{ij})_{j=1}^{n})_{i=1}^{n}$ with $d_{ij}\in\{0,1\}$ corresponds to the set of edges $\bm{r}$ $\bm{r}\equiv(r_{ij})_{(i,j)\in\mathcal{E}}$ with $r_{ij}\in\{0,1\}$ corresponds to the topological ordering $\bm{p}$ $\bm{p}\equiv(p_{i})_{i=1}^{I}$ with $p_{i}\in\{0,1\}$ is the set of primary variables $\bm{q}$ $\bm{q}\equiv(q_{i})_{i=1}^{J}$ with $q_{i}\in\{0,1\}$ is the set of auxiliary variables $\bm{u}$ $\bm{u}\equiv((u_{ij})_{j=1}^{\mu_{i}})_{i=1}^{n}$ with $u_{ij}\in\{0,1\}$ is the set of candidate parent subsets $\bm{v}$ $\bm{v}\equiv(v_{i})_{i=1}^{\nu}$ with $v_{i}\in\{0,1\}$ captures the search space of auxiliary variables $\bm{z}$ $\bm{z}\equiv(z_{i})_{i=1}^{n}$ with $z_{i}\in\{0,1\}$ is used to ensure that each vertex has exactly one parent set $\pi$ $\pi:2^{\{1,\cdots,I\}}\rightarrow\mathbb{R}\setminus\{0\}$ returns the coefficients in a multilinear polynomial $\bar{\pi}$ $\bar{\pi}:2^{\{1,\cdots,I\}}\times 2^{\{1,\cdots,I\}}\rightarrow\mathbb{R}$ returns the coefficients in a decomposed quadratization $\mathcal{F}$ $\mathcal{F}\subseteq 2^{\{1,\cdots,I\}}\setminus\{\emptyset\}$ is the terms in a multilinear polynomial $\mathcal{E}$ the set of edges on possible cycles given candidate parent sets $Q_{i}$ a subset of $2^{\{1,\cdots,I\}}\setminus\{\emptyset\}$ $W_{ij}$ a candidate parent set of $X_{i}$ $U_{ij}$ a candidate parent subset of $X_{i}$ $\mathcal{F}_{i}$ candidate parent sets of $X_{i}$ except for the empty set $\mathcal{X}_{i}$ the union of the candidate parent sets of $X_{i}$ $m_{i}$ the maximum parent set size of $X_{i}$ $S_{i}$ $S_{i}:2^{\mathcal{X}\setminus\{X_{i}\}}\rightarrow\mathbb{R}$ represents the local score function for a parent set of $X_{i}$ $\Pi_{i}$ the set of vertices that have edges directed towards $X_{i}$ $\pi_{i}$ $\pi_{i}(W)\equiv\sum_{W^{\prime}\subseteq W}(-1)^{1+|W\setminus W^{\prime}|}\log S_{i}(W^{\prime})$ $\bar{\pi}_{i}$ $\bar{\pi}_{i}(U,U^{\prime})\equiv(-1)^{1+|\{U\}\cup\{U^{\prime}\}|}(\log S_{i}(U\cup U^{\prime})-\log S_{i}(U)-\log S_{i}(U^{\prime})+\log S_{i}(\emptyset))$ $h$ $h:\{0,1\}^{J}\rightarrow[0,\infty)$ is a quadratic polynomial function $R$ $R(r_{ij},r_{jk},r_{ik})\equiv r_{ik}+r_{ij}r_{jk}-r_{ij}r_{ik}-r_{jk}r_{ik}$ $O$ $O(\bm{d})\equiv\sum_{1\leq i\leq n}\sum_{W\subseteq\mathcal{X}_{i}}\pi_{i}(W)\prod_{1\leq j\leq n;X_{j}\in W}d_{ij}$ $C$ $C(\bm{d},\bm{r})\equiv\sum_{(i,j),(j,k),(i,k)\in\mathcal{E}}\delta_{1}R(r_{ij},r_{jk},r_{ik})+\sum_{(i,j)\in\mathcal{E}}\delta_{2}\bigl{(}d_{ij}r_{ij}+d_{ji}(1-r_{ij})\bigr{)}$ $\bar{O}$ $\bar{O}(\bm{u},\bm{z})\equiv\sum_{1\leq i\leq n}\bigl{(}\sum_{1\leq j\leq k\leq\mu_{i}}\bar{\pi}_{i}(U_{ij},U_{ik})u_{ij}u_{ik}+\xi\bigl{(}z_{i}-z_{i}\sum_{1\leq j\leq\mu_{i}}u_{ij}+\sum_{1\leq j<k\leq\mu_{i}}u_{ij}u_{ik}\bigr{)}\bigr{)}$ $\bar{C}$ $\bar{C}(\bm{u},\bm{r})$ is defined as $C(\bm{d},\bm{r})$ where $d_{ij}$ is replaced by $\sum_{1\leq k\leq\mu_{i};X_{j}\in U_{ik}}u_{ik}$ $D$ $D(\bm{u})\equiv((\min\{1,\sum_{1\leq k\leq\mu_{i};X_{j}\in U_{ik}}u_{ik}\})_{j=1}^{n})_{i=1}^{n}$ $\bar{H}$ $\bar{H}(\bm{u},\bm{r},\bm{z})\equiv\bar{O}(\bm{u},\bm{z})+\bar{C}(\bm{u},\bm{r})$ $\bar{g}_{1}$ $\bar{g}_{1}(\bm{q})\equiv(\min\{1,\sum_{1\leq j\leq J;i\in Q_{j}}q_{j}\})_{i=1}^{I}$ $\bar{g}_{2}$ $\bar{g}_{2}(\bm{q})\equiv\sum_{1\leq i\leq j\leq J}\bar{\pi}(Q_{i},Q_{j})q_{i}q_{j}+h(\bm{q})$ $\xi_{0}$ $\xi_{0}\equiv-3\min_{1\leq i\leq n}\min_{0\leq j\leq k\leq\mu_{i}}\bar{\pi}_{i}(U_{ij},U_{ik})$ $\delta_{0}$ $\delta_{0}\equiv\max_{1\leq i\leq n,0\leq j\leq\lambda_{i}}(\log S_{i}(W_{ij})-\log S_{i}(\emptyset))$ $V_{\mathrm{close}}$ a subset of an optimal solution $V_{\mathrm{open}}$ $V_{\mathrm{close}}\cup V_{\mathrm{open}}$ contains an optimal solution $W_{\mathrm{open}}$ the terms that still cannot be ignored in $\mathcal{F}$
Appendix B Proof
B.1 Proof of Theorem 3.5
Initially, we introduce the following lemmas.
Lemma B.1.
Let $h_{0}$ be defined as
$h_{0}(\bm{q})\equiv\sum_{1\leq i\leq J}\sum_{\begin{subarray}{c}\emptyset\subset\mathcal{W}\subseteq\{1,\cdots,J\}\setminus\{i\}\\ Q_{i}\subseteq\bigcup_{j\in\mathcal{W}}Q_{j}\mathrm{~{}and~{}}Q_{i}\nsubseteq\bigcup_{j\in\mathcal{W}^{\prime}}Q_{j}\mathrm{~{}for~{}any~{}}\mathcal{W}^{\prime}\subset\mathcal{W}\end{subarray}}(1-q_{i})\prod_{j\in\mathcal{W}}q_{j}.$
Suppose that $\bm{p}=\bar{g}_{1}(\bm{q})$ holds. Then we have
$h_{0}(\bm{q})=0\Rightarrow\bm{q}=(\prod_{1\leq i\leq I;i\in Q_{j}}p_{i})_{j=1}^{J}\mathrm{~{}~{}for~{}any~{}}\bm{q}\in\{0,1\}^{J}.$
Proof.
From $\bm{p}=\bar{g}_{1}(\bm{q})$, if $p_{i}=1$, then there exists an index $j$ such that $i\in Q_{j}$ and $q_{j}=1$. Therefore, if $h_{0}(\bm{q})=0$, then $Q_{j}\subseteq\bigcup_{1\leq i\leq I;p_{i}=1}\{i\}\Rightarrow q_{j}=1$. Additionally, if $p_{i}=0$, then $i\in Q_{j}\Rightarrow q_{j}=0$. Consequently, this lemma holds. ∎
Lemma B.2.
Let $h$ be defined as
$h(\bm{q})\equiv\kappa\Bigl{(}\sum_{1\leq i\leq J}\sum_{\begin{subarray}{c}1\leq j<k\leq J\\ i\notin\{j,k\},Q_{j}\cup Q_{k}=Q_{i}\end{subarray}}(q_{j}q_{k}-2q_{j}q_{i}-2q_{k}q_{i}+3q_{i})+\sum_{\begin{subarray}{c}i,j\in\{1,\cdots,J\}\\ Q_{i}\subset Q_{j}\end{subarray}}(1-q_{i})q_{j}\Bigr{)},$
where $\kappa\in(0,\infty)$. Suppose that assumption 3.4 holds. Then we have
$h(\bm{q})=0\Rightarrow h_{0}(\bm{q})=0\mathrm{~{}~{}for~{}any~{}}\bm{q}\in\{0,1\}^{J}.$
Proof.
We prove $h_{0}(\bm{q})\neq 0\Rightarrow h(\bm{q})\neq 0$ for any $\bm{q}\in\{0,1\}^{J}$. If $h_{0}(\bm{q})\neq 0$, there exists at least one pair of an index $i$ and a set $\mathcal{W}$ such that $q_{i}=0$ and $\prod_{j\in\mathcal{W}}q_{j}=1$ in $h_{0}(\bm{q})$. For $|\mathcal{W}|=1$, the second term of $h(\bm{q})$ is positive. Consider the case where $|\mathcal{W}|\geq 2$. From assumption 3.4, there exists $(a,b)$ such that $Q_{a}\cup Q_{b}=Q_{i}$ and $i\notin\{a,b\}$. By repeating this split, we reach $Q_{c}$ such that $Q_{c}\subseteq Q_{j}$ for some $j\in\mathcal{W}$. Therefore, either the first or the second term of $h(\bm{q})$ is positive. Consequently, this lemma holds. ∎
Let $\bar{\pi}$ satisfy
$\sum_{1\leq i\leq j\leq J;Q_{i}\cup Q_{j}=V}\bar{\pi}(Q_{i},Q_{j})=\left\{\begin{array}[]{ll}\pi(V)&(V\in\mathcal{F})\\ 0&(V\notin\mathcal{F}).\end{array}\right.$ (22)
From lemma B.1 and lemma B.2, the following holds:
$h(\bm{q})=0\Rightarrow\bm{q}=(\prod_{1\leq i\leq I;i\in Q_{j}}p_{i})_{j=1}^{J}\mathrm{~{}~{}for~{}any~{}}\bm{q}\in\{0,1\}^{J}.$
To induce $h(\bm{q})=0$, we demonstrate a simple bound. Let $\kappa_{0}$ denote
$\kappa_{0}\equiv\max_{\bm{q}\in\{0,1\}^{J}}\sum_{1\leq i\leq j\leq J}\bar{\pi}(Q_{i},Q_{j})q_{i}q_{j}-\min_{\bm{q}\in\{0,1\}^{J}}\sum_{1\leq i\leq j\leq J}\bar{\pi}(Q_{i},Q_{j})q_{i}q_{j}.$
Suppose that $\kappa>\kappa_{0}$ holds. Then the following holds:
$f(\bm{p})=\min_{\bm{q}\in\{0,1\}^{J};\bm{p}=\bar{g}_{1}(\bm{q})}\bar{g}_{2}(\bm{q}).$
Hence, under assumption 3.3 and assumption 3.4, $(\bar{g}_{1},\bar{g}_{2})$ is a decomposed quadratization of $f$.
B.2 Proof of Lemma 3.6
Considering $\phi(1,\cdots,1,0)\neq 0$ and ${}_{K}\mathrm{C}_{2}+{}_{K}\mathrm{C}_{1}+{}_{K}\mathrm{C}_{0}<2^{K}$, no quadratic polynomial function $\phi:\{0,1\}^{K+1}\rightarrow\mathbb{R}$ satisfies eq. 11 for $K\in\{3,4,\cdots\}$.
B.3 Proof of Lemma 3.7
Let $\phi$ be
$\phi(q_{1},q_{2},q_{3})=b_{1}q_{1}+b_{2}q_{2}+b_{3}q_{3}+b_{12}q_{1}q_{2}+b_{13}q_{1}q_{3}+b_{23}q_{2}q_{3}.$
From eq. 11, $b_{1}=b_{2}=0$ and $b_{12}=\phi(1,1,0)\neq 0$. For $q_{3}=1$, the following holds:
$\phi(0,0,1)=b_{3},$
$\phi(1,0,1)=b_{1}+b_{3}+b_{13}=b_{3}+b_{13},$
$\phi(0,1,1)=b_{2}+b_{3}+b_{23}=b_{3}+b_{23},$
$\phi(1,1,1)=b_{1}+b_{2}+b_{3}+b_{12}+b_{13}+b_{23}=\phi(1,1,0)+b_{3}+b_{13}+b_{23}.$
Consequently, we have
$\phi(0,0,1)+\phi(1,1,1)-\phi(1,0,1)-\phi(0,1,1)=\phi(1,1,0)\neq 0.$
B.4 Proof of Lemma 4.3
For $u_{ab}=1$ and $\sum_{1\leq j\leq\mu_{a}}u_{aj}>2$, the following holds:
$\min_{\bm{z}\in\{0,1\}^{n}}\bar{O}(\bm{u},\bm{z})-\min_{\bm{z}\in\{0,1\}^{n}}\bar{O}(\bm{u}^{ab},\bm{z})\geq(-2+\sum_{1\leq j\leq\mu_{a}}u_{aj})\xi+\sum_{1\leq j\leq\mu_{a}}\bar{\pi}_{a}(U_{ab},U_{aj})u_{aj},$
where $\bm{u}^{ab}\equiv((u_{ij}^{ab})_{j=1}^{\mu_{i}})_{i=1}^{n}$ with $u_{ij}^{ab}=0$ if $(i,j)=(a,b)$; $u_{ij}^{ab}=u_{ij}$ otherwise. If $\xi>\xi_{0}$, then we have
$\min_{\bm{z}\in\{0,1\}^{n}}\bar{O}(\bm{u},\bm{z})>\min_{\bm{z}\in\{0,1\}^{n}}\bar{O}(\bm{u}^{ab},\bm{z}).$
This inequality implies that $\sum_{1\leq j\leq\mu_{i}}u_{ij}\leq 2$ holds for all $1\leq i\leq n$. Consequently, if $\xi>\xi_{0}$ holds and the state of $\bm{d}$ aligns with candidate parent sets, then $(D,\bar{O})$ is a decomposed quadratization of $O$.
B.5 Proof of Theorem 4.4
At first, we introduce the following lemmas.
Lemma B.3.
Let $T$ be defined as
$T(\bm{r})\equiv\sum_{(i,j),(j,k),(i,k)\in\mathcal{E}}R(r_{ij},r_{jk},r_{ik}).$
Suppose that $T(\bm{r})>0$ holds. Then there exists at least one integer pair $(a,b)$ such that $T(\bm{r})>T(\bm{r}^{ab})$, where $\bm{r}^{ab}\equiv(r_{ij}^{ab})_{(i,j)\in\mathcal{E}}$ with $r_{ij}^{ab}=1-r_{ij}$ if $(i,j)=(a,b)$; $r_{ij}^{ab}=r_{ij}$ otherwise.
Proof.
Assume $R(r_{ij},r_{jk},r_{ik})>0$. For all $l$ satisfying $k<l\leq n$, the following holds:
$R(r_{ij},r_{jl},r_{il})-R(1-r_{ij},r_{jl},r_{il})+R(r_{jk},r_{kl},r_{jl})-R(1-r_{jk},r_{kl},r_{jl})+R(r_{ik},r_{kl},r_{il})-R(1-r_{ik},r_{kl},r_{il})$
$=(r_{il}+r_{ij}r_{jl}-r_{ij}r_{il}-r_{jl}r_{il})-(r_{jl}+r_{ij}r_{il}-r_{ij}r_{jl}-r_{jl}r_{il})$
$~{}~{}~{}+(r_{jl}+r_{jk}r_{kl}-r_{jk}r_{jl}-r_{kl}r_{jl})-(r_{kl}+r_{jk}r_{jl}-r_{jk}r_{kl}-r_{kl}r_{jl})$
$~{}~{}~{}+(r_{il}+r_{ik}r_{kl}-r_{ik}r_{il}-r_{kl}r_{il})-(r_{kl}+r_{ik}r_{il}-r_{ik}r_{kl}-r_{kl}r_{il})$
$=2r_{il}(1-r_{ij}-r_{ik})+2r_{jl}(r_{ij}-r_{jk})+2r_{kl}(r_{jk}+r_{ik}-1)$
$=0,$
where $(r_{ij},r_{jk},r_{ik})$ is $(1,1,0)$ or $(0,0,1)$. For all $l$ satisfying $j<l<k$, the following holds:
$R(r_{ij},r_{jl},r_{il})-R(1-r_{ij},r_{jl},r_{il})+R(r_{jl},r_{lk},r_{jk})-R(r_{jl},r_{lk},1-r_{jk})+R(r_{il},r_{lk},r_{ik})-R(r_{il},r_{lk},1-r_{ik})$
$=(r_{il}+r_{ij}r_{jl}-r_{ij}r_{il}-r_{jl}r_{il})-(r_{jl}+r_{ij}r_{il}-r_{ij}r_{jl}-r_{jl}r_{il})$
$~{}~{}~{}+(r_{jk}+r_{jl}r_{lk}-r_{jl}r_{jk}-r_{lk}r_{jk})-(1-r_{jk}-r_{lk}-r_{jl}+r_{jl}r_{lk}+r_{jl}r_{jk}+r_{lk}r_{jk})$
$~{}~{}~{}+(r_{ik}+r_{il}r_{lk}-r_{il}r_{ik}-r_{lk}r_{ik})-(1-r_{ik}-r_{il}-r_{lk}+r_{il}r_{lk}+r_{il}r_{ik}+r_{lk}r_{ik})$
$=2r_{il}(1-r_{ik}-r_{ij})+2r_{lk}(1-r_{ik}-r_{jk})+2r_{jl}(r_{ij}-r_{jk})+2(r_{ik}+r_{jk}-1)$
$=0.$
For all $l$ satisfying $i<l<j$, the following holds:
$R(r_{il},r_{lj},r_{ij})-R(r_{il},r_{lj},1-r_{ij})+R(r_{lj},r_{jk},r_{lk})-R(r_{lj},1-r_{jk},r_{lk})+R(r_{il},r_{lk},r_{ik})-R(r_{il},r_{lk},1-r_{ik})$
$=(r_{ij}+r_{il}r_{lj}-r_{il}r_{ij}-r_{lj}r_{ij})-(1-r_{ij}-r_{il}-r_{lj}+r_{il}r_{lj}+r_{il}r_{ij}+r_{lj}r_{ij})$
$~{}~{}~{}+(r_{lk}+r_{lj}r_{jk}-r_{lj}r_{lk}-r_{jk}r_{lk})-(r_{lj}+r_{jk}r_{lk}-r_{lj}r_{jk}-r_{lj}r_{lk})$
$~{}~{}~{}+(r_{ik}+r_{il}r_{lk}-r_{il}r_{ik}-r_{lk}r_{ik})-(1-r_{ik}-r_{il}-r_{lk}+r_{il}r_{lk}+r_{il}r_{ik}+r_{lk}r_{ik})$
$=2r_{il}(1-r_{ik}-r_{ij})+2r_{lk}(1-r_{ik}-r_{jk})+2(r_{ij}+r_{ik}-1)$
$=0.$
For all $l$ satisfying $1\leq l<i$, the following holds:
$R(r_{li},r_{ij},r_{lj})-R(r_{li},1-r_{ij},r_{lj})+R(r_{lj},r_{jk},r_{lk})-R(r_{lj},1-r_{jk},r_{lk})+R(r_{li},r_{ik},r_{lk})-R(r_{li},1-r_{ik},r_{ik})$
$=(r_{lj}+r_{li}r_{ij}-r_{li}r_{lj}-r_{ij}r_{lj})-(r_{li}+r_{ij}r_{lj}-r_{li}r_{ij}-r_{li}r_{lj})$
$~{}~{}~{}+(r_{lk}+r_{lj}r_{jk}-r_{lj}r_{lk}-r_{jk}r_{lk})-(r_{lj}+r_{jk}r_{lk}-r_{lj}r_{jk}-r_{lj}r_{lk})$
$~{}~{}~{}+(r_{lk}+r_{li}r_{ik}-r_{li}r_{lk}-r_{ik}r_{lk})-(r_{li}+r_{ik}r_{lk}-r_{li}r_{ik}-r_{li}r_{lk})$
$=2r_{lk}(1-r_{ik}-r_{jk})+2r_{li}(r_{ij}+r_{ik}-1)+2r_{lj}(r_{jk}-r_{ij})$
$=0.$
From $R(r_{ij},r_{jk},r_{ik})>0$, the following holds:
$R(r_{ij},r_{jk},r_{ik})-R(1-r_{ij},r_{jk},r_{ik})+R(r_{ij},r_{jk},r_{ik})-R(r_{ij},1-r_{jk},r_{ik})+R(r_{ij},r_{jk},r_{ik})-R(r_{ij},r_{jk},1-r_{ik})=3.$
Using these results, we have
$T(\bm{r})-T(\bm{r}^{ij})+T(\bm{r})-T(\bm{r}^{jk})+T(\bm{r})-T(\bm{r}^{ik})=3>0.$
Consequently, $T(\bm{r})>T(\bm{r}^{ab})$ holds for at least one index pair $(a,b)\in\{(i,j),(j,k),(i,k)\}$. ∎
Let $\bm{u}^{aef}\equiv((u_{ij}^{aef})_{j=1}^{\mu_{i}})_{i=1}^{n}$ satisfy
$u_{ae}^{aef}=u_{af}^{aef}=1,~{}~{}u_{aj}^{aef}=0\mathrm{~{}for~{}}j\notin\{e,f\},\mathrm{~{}and~{}}u_{ij}^{aef}=u_{ij}\mathrm{~{}for~{}}i\neq a.$
Assume the following conditions:
$u_{a0}\in\{0,1\},~{}~{}(a,b)\in\mathcal{E},~{}~{}0\leq c<d\leq\mu_{a},~{}~{}0\leq e\leq f\leq\mu_{a},~{}~{}u_{ac}=u_{ad}=1,$
$u_{a0}+\cdots+u_{a\mu_{a}}=2,~{}~{}X_{b}\in(U_{ac}\cup U_{ad})\setminus(U_{ae}\cup U_{af}),~{}~{}(U_{ae}\cup U_{af})\subset(U_{ac}\cup U_{ad}).$
We find the range of penalty coefficients so that the difference in the return value of the objective function is negative when the input state changes to the desired state. From the discussion in section B.4, we consider the case where $\sum_{1\leq j\leq\mu_{i}}u_{ij}\leq 2$ for all $1\leq i\leq n$. For $\sum_{1\leq j\leq\mu_{b};X_{a}\in U_{bj}}u_{bj}\geq 1$, the following holds:
$\min_{\bm{z}\in\{0,1\}^{n}}\bar{H}(\bm{u},\bm{r},\bm{z})-\min_{\bm{z}\in\{0,1\}^{n}}\bar{H}(\bm{u}^{aef},\bm{r},\bm{z})\geq\delta_{2}-\log S_{a}(U_{ac}\cup U_{ad})+\log S_{a}(U_{ae}\cup U_{af}).$
For $\sum_{1\leq j\leq\mu_{b};X_{a}\in U_{bj}}u_{bj}=0$ and $r_{ab}=1$, the following holds:
$\min_{\bm{z}\in\{0,1\}^{n}}\bar{H}(\bm{u},\bm{r},\bm{z})-\min_{\bm{z}\in\{0,1\}^{n}}\bar{H}(\bm{u},\bm{r}^{ab},\bm{z})\geq\delta_{2}-(n-2)\delta_{1}.$
From lemma B.3, we could repeat reversing one element from $\bm{r}$ with decreasing $T(\bm{r})$ until $T(\bm{r})=0$. For $T(\bm{r})>T(\bm{r}^{ab})$, $r_{ab}=0$, and $\sum_{1\leq j\leq\mu_{b};X_{a}\in U_{bj}}u_{bj}=0$, the following holds:
$\min_{\bm{z}\in\{0,1\}^{n}}\bar{H}(\bm{u},\bm{r},\bm{z})-\min_{\bm{z}\in\{0,1\}^{n}}\bar{H}(\bm{u}^{aef},\bm{r}^{ab},\bm{z})\geq\delta_{1}-\log S_{a}(U_{ac}\cup U_{ad})+\log S_{a}(U_{ae}\cup U_{af}).$
Here the following holds:
$\delta_{0}\geq\max_{1\leq a\leq n}\max_{1\leq b\leq n}\max_{0\leq c<d\leq\mu_{a}}\min_{\begin{subarray}{c}0\leq e\leq f\leq\mu_{a}\\ X_{b}\in(U_{ac}\cup U_{ad})\setminus(U_{ae}\cup U_{af}),(U_{ae}\cup U_{af})\subset(U_{ac}\cup U_{ad})\end{subarray}}\log\frac{S_{a}(U_{ac}\cup U_{ad})}{S_{a}(U_{ae}\cup U_{af})}.$
From the above discussion, if $\xi>\xi_{0}$ and $\delta_{2}>(n-2)\delta_{1}>(n-2)\delta_{0}$ hold, then the following holds:
$\bar{C}(\bm{u_{*}},\bm{r_{*}})=0\mathrm{~{}~{}for~{}any~{}}\bar{H}(\bm{u_{*}},\bm{r_{*}},\bm{z_{*}})=\min_{\bm{u},\bm{r},\bm{z}}\bar{H}(\bm{u},\bm{r},\bm{z}).$
Appendix C Integer Linear Programming for QUBO Formulation
C.1 Formulation for Reducing Binary Variables
For any $1\leq i\leq\nu$, define
$\mathcal{W}_{i}\equiv\{\mathcal{W}\subseteq V_{\mathrm{close}}\cup V_{\mathrm{open}}\mid|\mathcal{W}|\geq 2,V_{i}\subseteq\bigcup_{V\in\mathcal{W}}V,\mathrm{~{}and~{}}V_{i}\nsubseteq\bigcup_{V\in\mathcal{W}^{\prime}}V\mathrm{~{}~{}for~{}any~{}}\mathcal{W}^{\prime}\subset\mathcal{W}\}.$
The search for an optimal solution can be formulated as an integer linear programming problem to minimize $\sum_{1\leq i\leq\nu}v_{i}$ under the following constraints:
$\sum_{\begin{subarray}{c}1\leq i\leq\nu\\ W\in\bigcup_{V\in\{\emptyset\}\cup V_{\mathrm{close}}}\{V_{i}\cup V\}\end{subarray}}v_{i}+\sum_{\begin{subarray}{c}1\leq i<j\leq\nu\\ W\notin\{V_{i},V_{j}\},V_{i}\cup V_{j}=W\end{subarray}}v_{i}v_{j}\geq 1\mathrm{~{}~{}for~{}all~{}}W\in W_{\mathrm{open}},$
$v_{i}\prod_{\begin{subarray}{c}1\leq j\leq\nu\\ V_{j}\in\mathcal{W}\end{subarray}}v_{j}\leq\sum_{\begin{subarray}{c}V,V^{\prime}\in V_{\mathrm{close}}\\ V\cup V^{\prime}=V_{i}\end{subarray}}1+\sum_{\begin{subarray}{c}j\in\{1,\cdots,\nu\}\setminus\{i\}\\ V_{i}\in\bigcup_{V\in V_{\mathrm{close}}}\{V_{j}\cup V\}\end{subarray}}v_{j}+\sum_{\begin{subarray}{c}1\leq j<k\leq\nu\\ i\notin\{j,k\},V_{j}\cup V_{k}=V_{i}\end{subarray}}v_{j}v_{k}$
$\mathrm{~{}~{}for~{}all~{}}1\leq i\leq\nu\mathrm{~{}and~{}}\mathcal{W}\in\mathcal{W}_{i}.$
Replacing $v_{i}v_{j}$ with $\rho_{ij}\in\{0,1\}$, we transform a higher degree terms into a linear as
$\sum_{\begin{subarray}{c}1\leq i\leq\nu\\ W\in\bigcup_{V\in\{\emptyset\}\cup V_{\mathrm{close}}}\{V_{i}\cup V\}\end{subarray}}v_{i}+\sum_{\begin{subarray}{c}1\leq i<j\leq\nu\\ W\notin\{V_{i},V_{j}\},V_{i}\cup V_{j}=W\end{subarray}}\rho_{ij}\geq 1\mathrm{~{}~{}for~{}all~{}}W\in W_{\mathrm{open}},$
$-|\mathcal{W}\setminus V_{\mathrm{close}}|+v_{i}+\sum_{\begin{subarray}{c}1\leq j\leq\nu\\ V_{j}\in\mathcal{W}\end{subarray}}v_{j}\leq\sum_{\begin{subarray}{c}V,V^{\prime}\in V_{\mathrm{close}}\\ V\cup V^{\prime}=V_{i}\end{subarray}}1+\sum_{\begin{subarray}{c}j\in\{1,\cdots,\nu\}\setminus\{i\}\\ V_{i}\in\bigcup_{V\in V_{\mathrm{close}}}\{V_{j}\cup V\}\end{subarray}}v_{j}+\sum_{\begin{subarray}{c}1\leq j<k\leq\nu\\ i\notin\{j,k\},V_{j}\cup V_{k}=V_{i}\end{subarray}}\rho_{jk}$
$\mathrm{~{}for~{}all~{}}1\leq i\leq\nu\mathrm{~{}and~{}}\mathcal{W}\in\mathcal{W}_{i},$
$2\rho_{ij}\leq v_{i}+v_{j}\leq 1+\rho_{ij}\mathrm{~{}for~{}all~{}}1\leq i<j\leq\nu.$
To find optimal candidate parent subsets, we ignore the latter constraint.
C.2 Formulation for Term Reduction
Consider the formulation shown in section B.1. Let $\eta_{ij}\in\{0,1\}$ represent the existence of a term $v_{i}v_{j}$. The number of terms is reduced by solving an integer programming problem that minimizes $\sum_{1\leq i\leq j\leq\nu+\nu^{\prime}}\eta_{ij}$ subject to the following constraints:
$\eta_{ij}\leq v_{i}v_{j}\mathrm{~{}~{}for~{}all~{}}1\leq i\leq j\leq\nu+\nu^{\prime},$
$v_{i}=1\mathrm{~{}~{}for~{}all~{}}\nu+1\leq i\leq\nu+\nu^{\prime},$
$\eta_{ij}=1\mathrm{~{}~{}for~{}all~{}}\nu+1\leq i\leq j\leq\nu+\nu^{\prime}\mathrm{~{}and~{}}V_{i}\cup V_{j}\in\mathcal{F},$
$v_{i}v_{j}\leq\frac{1}{2}(\eta_{jj}+\eta_{\min\{i,j\}\max\{i,j\}})\mathrm{~{}~{}for~{}all~{}}i,j\in\{1,\cdots,\nu+\nu^{\prime}\}\mathrm{~{}and~{}}V_{i}\subset V_{j},$
$-2+v_{i}+v_{j}+v_{k}\leq\frac{1}{4}(\eta_{ii}+\eta_{\min\{i,j\}\max\{i,j\}}+\eta_{\min\{i,k\}\max\{i,k\}}+\eta_{jk})$
$\mathrm{~{}~{}for~{}all~{}}1\leq i\leq\nu+\nu^{\prime},1\leq j<k\leq\nu+\nu^{\prime},i\notin\{j,k\},\mathrm{~{}and~{}}V_{j}\cup V_{k}=V_{i},$
$\sum_{\begin{subarray}{c}1\leq i\leq j\leq\nu+\nu^{\prime}\\ V_{i}\cup V_{j}=W\end{subarray}}\eta_{ij}\geq 1\mathrm{~{}~{}for~{}all~{}}W\in\mathcal{F},$
$-|\mathcal{W}|+v_{i}+\sum_{\begin{subarray}{c}1\leq j\leq\nu+\nu^{\prime}\\ V_{j}\in\mathcal{W}\end{subarray}}v_{j}\leq\sum_{\begin{subarray}{c}1\leq j<k\leq\nu+\nu^{\prime}\\ i\notin\{j,k\},V_{j}\cup V_{k}=V_{i}\end{subarray}}v_{j}v_{k}\mathrm{~{}~{}for~{}all~{}}1\leq i\leq\nu\mathrm{~{}and~{}}\mathcal{W}\in\mathcal{W}_{i},$
where $(v_{i})_{i=\nu+1}^{\nu+\nu^{\prime}}\in\{0,1\}^{\nu^{\prime}}$ and $\{V_{i}\}_{i=\nu+1}^{\nu+\nu^{\prime}}\equiv V_{\mathrm{close}}$. As described in section C.1, we transform quadratic terms into linear ones by replacing $v_{i}v_{j}$ with $\rho_{ij}$.
Appendix D Experiment
D.1 Bayesian Dirichlet Equivalence Uniform Score
For the Bayesian Dirichlet equivalence uniform (BDeu) score, define
$S_{i}(\Pi)\equiv\prod_{1\leq j\leq\beta_{i}}\frac{\Gamma(\alpha_{ij})}{\Gamma(N_{ij}+\alpha_{ij})}\prod_{1\leq k\leq\gamma_{i}}\frac{\Gamma(N_{ijk}+\alpha_{ijk})}{\Gamma(\alpha_{ijk})}\mathrm{~{}~{}for~{}any~{}}1\leq i\leq n\mathrm{~{}and~{}}|\Pi|\leq m_{i},$
where $N_{ij}\equiv\sum_{1\leq k\leq\gamma_{i}}N_{ijk}$, $N_{ijk}$ is the number of cases of $\Pi$ in its $j$-th state and $X_{i}$ in its $k$-th state, $\alpha_{ij}\equiv\sum_{1\leq k\leq\gamma_{i}}\alpha_{ijk}$, $\alpha_{ijk}\equiv\frac{\alpha}{\beta_{i}\gamma_{i}}$, $\beta_{i}$ is the number of joint states of $\Pi$, $\gamma_{i}$ is the number of states of $X_{i}$, and $\alpha\in(0,\infty)$ is the equivalent sample size.
D.2 Identification of Candidate Parent Sets
[] Instance $m$ Number Time [s] alarm $4$ $2291\pm 210$ $1700\pm 11$ barley $4$ $355\pm 12$ $235015\pm 95243$ hailfinder $4$ $692\pm 16$ $90247\pm 3334$ hepar2 $4$ $689\pm 59$ $40216\pm 534$ chess $2$ $32570\pm 1686$ $59\pm 0$ chess $3$ $236285\pm 15636$ $5463\pm 284$ win95pts $2$ $9231\pm 665$ $60\pm 3$ win95pts $3$ $43657\pm 3850$ $2404\pm 51$ pathfinder $2$ $50962\pm 3254$ $320\pm 5$ pathfinder $3$ $398131\pm 47889$ $44991\pm 3426$ mushroom $2$ $273526\pm 12647$ $689\pm 26$ connect $2$ $76996\pm 4381$ $331\pm 7$ connect $3$ $710684\pm 73964$ $83361\pm 6624$ munin1 $2$ $135570\pm 6765$ $2466\pm 29$ andes $2$ $7424\pm 219$ $852\pm 52$ andes $3$ $14195\pm 556$ $104683\pm 3855$
Table 4: Number of candidate parent sets and execution time for their identification. The mean and standard deviation of $5$ trials are presented. Empty sets were excluded from the candidate parent sets.
D.3 Classical Simulated Annealing
Algorithm 3 Classical Simulated Annealing
1:  while the current time is within the time limit, do
2:     $\Delta\leftarrow$ the incremental return value of $\bar{H}$ when filipping a bit randomly.
3:     if a random number between 00 and $1$ is less than $\exp(-\Delta/T_{0}^{1-\mathrm{current~{}time}/\mathrm{time~{}limit}})$then
4:        accept the bit flip.
5:     end if
6:  end while
D.4 Binary Quadratic Programming
The binary quadratic programming problem for learning a Bayesian network can be defined by
$\mathrm{minimize~{}}\sum_{1\leq i\leq n}\sum_{1\leq j\leq k\leq\mu_{i}}\bar{\pi}_{i}(U_{ij},U_{ik})u_{ij}u_{ik}+\bar{C}(\bm{u},\bm{r})\mathrm{~{}~{}~{}~{}subject~{}to~{}}\sum_{1\leq j\leq\mu_{i}}u_{ij}\leq 2\mathrm{~{}~{}for~{}all~{}}1\leq i\leq n.$
Note that we can express the DAG constraint using linear inequalities. The linear ordering of vertices is
$0\leq r_{ij}+r_{jk}-r_{ik}\leq 1\mathrm{~{}~{}for~{}all~{}}(i,j),(j,k),(i,k)\in\mathcal{E}.$
The consistency of edges and ordering is
$\sum_{\begin{subarray}{c}1\leq k\leq\mu_{i}\\ X_{j}\in U_{k}\end{subarray}}u_{ik}\leq 2-2r_{ij}\mathrm{~{}~{}and~{}~{}}\sum_{\begin{subarray}{c}1\leq k\leq\mu_{j}\\ X_{i}\in U_{k}\end{subarray}}u_{jk}\leq 2r_{ij}\mathrm{~{}~{}~{}for~{}all~{}}(i,j)\in\mathcal{E}.$
D.5 Identification of Candidate Parent Subsets
[] Instance $m$ Time [s] alarm $4$ $1\pm 0,50\pm 28,10\pm 1$ barley $4$ $0\pm 0,1\pm 0,10\pm 2$ hailfinder $4$ $0\pm 0,0\pm 0,4\pm 1$ hepar2 $4$ $0\pm 0,0\pm 0,2\pm 0$ chess $2$ $6\pm 0,0\pm 0,83\pm 3$ chess $3$ $639\pm 58,3150\pm 167,3480\pm 315$ win95pts $2$ $2\pm 2,0\pm 0,31\pm 1$ win95pts $3$ $49\pm 8,530\pm 73,457\pm 66$ pathfinder $2$ $18\pm 2,0\pm 0,241\pm 29$ pathfinder $3$ $3844\pm 750,2132\pm 129,24731\pm 3458$ mushroom $2$ $277\pm 19,1\pm 0,403\pm 14$ connect $2$ $26\pm 3,0\pm 0,203\pm 15$ connect $3$ $6219\pm 1396,4568\pm 177,19595\pm 2743$ munin1 $2$ $78\pm 6,0\pm 0,875\pm 71$ andes $2$ $1\pm 0,0\pm 0,247\pm 37$ andes $3$ $4\pm 0,7\pm 1,493\pm 131$
Table 5: Execution time for IP only. Displays are the same as table 2.
D.6 Minor Embedding
Annealing machines are classified into two types: nearest-neighbor and fully-coupled. A fully-coupled annealing machine allows coupling between arbitrary vertices, whereas coupling in a nearest-neighbor annealing machine is limited to adjacent vertices only. Nearest-neighbor annealing machines require additional bits to embed a highly dense logical topology into a physical one (Choi 2008, 2010). The number of additional bits for the embedding depends on the design of the hardware graphs, and embedding problems are NP-hard (Eppstein 2009).

### v8

@@
 2014).
-These instances can be encoded on the circuit of the Fujitsu Digital Annealer, a fully-coupled annealing machine with $100$K bit capacity (Nakayama et al.
+These instances can be encoded on the circuit of the Fujitsu Digital Annealer, a fully-coupled annealing machine with 100100K bit capacity (Nakayama et al.
 2021).
@@
 Example 3.9.
-Let =$\mathcal{F}=\{\{1,2,3,4\},\{1,2,3\},\{4\},\{3,4\}\}$.
+Let $\mathcal{F}=\{\{1,2,3,4\},\{1,2,3\},\{4\},\{3,4\}\}$.
 A decomposed quadratization requires $3$ auxiliary variables: $Q_{1}=\{1,2,3\},Q_{2}=\{4\}$, and $Q_{3}=\{3\}$.
@@
 Algorithm 1 Search Space Reduction
-1:  $W_{\mathrm{open}}\leftarrow\mathcal{F},V_{\mathrm{close}}\leftarrow\emptyset,V_{\mathrm{open}}\leftarrow\bigcup_{V\in\mathcal{F}}2^{V}\setminus\{\emptyset\}$.
+[…]
 2: Merge $\{V\in V_{\mathrm{open}}\mid$ for some $W\in W_{\mathrm{open}}$, $V$ is only an element in $V_{\mathrm{open}}$ such that $W\in\bigcup_{V^{\prime}\in\{\emptyset\}\cup V_{\mathrm{close}}}\{V\cup V^{\prime}\}\}$ into $V_{\mathrm{close}}$.
@@
 An undirected edge between two binary variables represents that the objective function contains their product term.
-Gray signifies that the state of a binary variable is 00 and white is $1$.
+Gray signifies that the state of a binary variable is 0 and white is $1$.
 Each color corresponds to the edge in the DAG.
@@
 Algorithm 2 Candidate Parent Subset Reduction
-1:  $W_{\mathrm{open}}\leftarrow\mathcal{F}_{i},V_{\mathrm{close}}\leftarrow\emptyset,V_{\mathrm{open}}\leftarrow\bigcup_{V\in\mathcal{F}_{i}}2^{V}\setminus\{\emptyset\}$.
+[…]
 2:while $V_{\mathrm{close}}$ has changed since the last evaluation, do
 3:  Exclude $\{V\in V_{\mathrm{open}}\mid V\subseteq W$ for at most one element $W\in W_{\mathrm{open}}\}$ from $V_{\mathrm{open}}$.
-4:     Execute the row from $2$ to $4$ in algorithm 1.
+4:  Execute the row from $2$ to 44 in algorithm 1.
 5:  Merge $W_{\mathrm{open}}\setminus\bigcup_{V,V^{\prime}\in V_{\mathrm{close}}\cup V_{\mathrm{open}}}\{V\cup V^{\prime}\}$ into $V_{\mathrm{close}}$, and exclude it from $W_{\mathrm{open}}$.
@@
 Additionally, we can exclude the constraint that captures assumption 3.4 from the integer linear programming problem.
-[] Instance $m$ O’Gorman et al.
-IP only IP + algorithm 2 Number Time [s] Number Time [s] Number Time [s] alarm $4$ $3982\pm 346$ $272\pm 61$ $\bm{1373\pm 46}$ $61\pm 29$ $\bm{1373\pm 46}$ $\bm{14\pm 2}$ barley $4$ $1826\pm 279$ $271\pm 149$ $\bm{995\pm 143}$ $\bm{11\pm 2}$ $\bm{995\pm 143}$ $\bm{11\pm 2}$ hailfinder $4$ $2374\pm 31$ $52\pm 24$ $\bm{1913\pm 6}$ $\bm{4\pm 1}$ $\bm{1913\pm 6}$ $\bm{4\pm 1}$ hepar2 $4$ $3629\pm 423$ $51\pm 16$ $\bm{2616\pm 133}$ $3\pm 0$ $\bm{2616\pm 133}$ $\bm{2\pm 0}$ chess $2$ $7040\pm 167$ $\bm{63\pm 2}$ $\bm{6964\pm 165}$ $89\pm 4$ $\bm{6964\pm 165}$ $103\pm 4$ chess $3$ $80116\pm 3014$ $\bm{3229\pm 151}$ $26355\pm 1403$ $7268\pm 492$ $\bm{23662\pm 1103}$ $5567\pm 376$ win95pts $2$ $5150\pm 39$ $\bm{27\pm 2}$ $\bm{5028\pm 42}$ $33\pm 3$ $\bm{5028\pm 42}$ $37\pm 4$ win95pts $3$ $37705\pm 1935$ $1223\pm 90$ $9220\pm 398$ $1036\pm 117$ $\bm{9025\pm 333}$ $\bm{890\pm 105}$ pathfinder $2$ $10831\pm 382$ $\bm{178\pm 8}$ $\bm{10725\pm 379}$ $259\pm 30$ $\bm{10725\pm 379}$ $284\pm 20$ pathfinder $3$ $107725\pm 5329$ $\bm{11695\pm 834}$ $38245\pm 3242$ $30707\pm 4322$ $\bm{34448\pm 3016}$ $21364\pm 3773$ mushroom $2$ $18050\pm 508$ $\bm{318\pm 13}$ $\bm{17930\pm 506}$ $681\pm 32$ $\bm{17930\pm 506}$ $513\pm 23$ connect $2$ $15155\pm 334$ $\bm{137\pm 9}$ $\bm{14978\pm 340}$ $228\pm 17$ $\bm{14978\pm 340}$ $240\pm 12$ connect $3$ $179788\pm 6242$ $\bm{9897\pm 602}$ $64533\pm 4352$ $30381\pm 4267$ $\bm{58247\pm 4053}$ $20106\pm 2661$ munin1 $2$ $25180\pm 266$ $\bm{699\pm 20}$ $\bm{25001\pm 268}$ $953\pm 75$ $\bm{25001\pm 268}$ $1079\pm 67$ andes $2$ $29502\pm 171$ $\bm{40\pm 1}$ $\bm{28810\pm 182}$ $249\pm 37$ $\bm{28810\pm 182}$ $250\pm 35$ andes $3$ $92838\pm 3735$ $2182\pm 289$ $\bm{31162\pm 136}$ $504\pm 132$ $\bm{31162\pm 136}$ $\bm{217\pm 8}$
+[…]
 Table 1: Number of binary variables and execution time for the QUBO formulation.
-The mean and standard deviation of $5$ trials are presented.
+The mean and standard deviation of 55 trials are presented.
 All the candidate parent subsets not proven optimal within $60$ [s] were feasible solutions.
@@
 The comparison between IP + algorithm 2 and IP only in table 1 demonstrates that optimizing candidate parent subsets with algorithm 2 is advantageous in terms of both the number of binary variables and execution time when the parent set size is larger.
-The quadratization for some instances with $m=3$ required binary variables over the $100$K bit capacity of the Digital Annealer.
+The quadratization for some instances with $m=3$ required binary variables over the 100100K bit capacity of the Digital Annealer.
 The decomposed quadratization in our approach significantly reduced the required number of binary variables compared to the quadratization approach.
@@
 The Digital Annealer effectively highlighted the strengths of our formulation compared to the results from SAQ and GOQ using the same QUBO formulation.
-[] Instance $m$ Time [s] alarm $4$ $1\pm 0,3\pm 1,10\pm 1$ barley $4$ $0\pm 0,0\pm 0,11\pm 2$ hailfinder $4$ $0\pm 0,0\pm 0,4\pm 1$ hepar2 $4$ $0\pm 0,0\pm 0,2\pm 0$ chess $2$ $2\pm 0,0\pm 0,100\pm 4$ chess $3$ $406\pm 49,2554\pm 199,2608\pm 195$ win95pts $2$ $1\pm 0,0\pm 0,36\pm 4$ win95pts $3$ $45\pm 5,398\pm 82,446\pm 54$ pathfinder $2$ $3\pm 0,0\pm 0,281\pm 19$ pathfinder $3$ $1407\pm 272,1679\pm 184,18278\pm 3349$ mushroom $2$ $23\pm 1,0\pm 0,490\pm 22$ connect $2$ $5\pm 0,0\pm 0,235\pm 12$ connect $3$ $2472\pm 436,4129\pm 148,13505\pm 2114$ munin1 $2$ $12\pm 1,0\pm 0,1067\pm 67$ andes $2$ $1\pm 0,0\pm 0,249\pm 35$ andes $3$ $6\pm 1,2\pm 1,209\pm 8$
+[] Instance $m$ Time [s] alarm 44 $1\pm 0,3\pm 1,10\pm 1$ barley 44 $0\pm 0,0\pm 0,11\pm 2$ hailfinder 44 $0\pm 0,0\pm 0,4\pm 1$ hepar2 44 $0\pm 0,0\pm 0,2\pm 0$ chess 22 $2\pm 0,0\pm 0,100\pm 4$ chess 33 $406\pm 49,2554\pm 199,2608\pm 195$ win95pts 22 $1\pm 0,0\pm 0,36\pm 4$ win95pts 33 $45\pm 5,398\pm 82,446\pm 54$ pathfinder 22 $3\pm 0,0\pm 0,281\pm 19$ pathfinder 33 $1407\pm 272,1679\pm 184,18278\pm 3349$ mushroom 22 $23\pm 1,0\pm 0,490\pm 22$ connect 22 $5\pm 0,0\pm 0,235\pm 12$ connect 33 $2472\pm 436,4129\pm 148,13505\pm 2114$ munin1 22 $12\pm 1,0\pm 0,1067\pm 67$ andes 22 $1\pm 0,0\pm 0,249\pm 35$ andes 33 $6\pm 1,2\pm 1,209\pm 8$
 Table 2: Execution time for IP + algorithm 2.
@@
 Note that the execution time for IP only is shown in section D.5.
-[] Instance $m$ DAQ GOQ HCS SAO GOB alarm $4$ 00 00 00 00 $\bm{5}$ barley $4$ 00 00 00 00 $\bm{5}$ hailfinder $4$ 00 00 00 00 $\bm{5}$ hepar2 $4$ 00 00 00 00 $\bm{5}$ chess $2$ $\bm{5}$ 00 00 00 00 chess $3$ $\bm{5}$ 00 00 00 00 win95pts $2$ $1$ 00 00 00 $\bm{4}$ win95pts $3$ $2$ 00 00 $\bm{3}$ 00 pathfinder $2$ $\bm{4}$ 00 00 $1$ 00 pathfinder $3$ $\bm{5}$ 00 00 00 00 mushroom $2$ $\bm{5}$ 00 00 00 00 connect $2$ $\bm{5}$ 00 00 00 00 connect $3$ $\bm{2}$ $1$ $1$ $1$ 00 munin1 $2$ $\bm{5}$ 00 00 00 00 andes $2$ $\bm{5}$ 00 00 00 00 andes $3$ $\bm{5}$ 00 00 00 00
+[] Instance $m$ DAQ GOQ HCS SAO GOB alarm 44 0 0 0 0 $\bm{5}$ barley 44 0 0 0 0 $\bm{5}$ hailfinder 44 0 0 0 0 $\bm{5}$ hepar2 44 0 0 0 0 $\bm{5}$ chess 22 $\bm{5}$ 0 0 0 0 chess 33 $\bm{5}$ 0 0 0 0 win95pts 22 11 0 0 0 $\bm{4}$ win95pts 33 22 0 0 $\bm{3}$ 0 pathfinder 22 $\bm{4}$ 0 0 11 0 pathfinder 33 $\bm{5}$ 0 0 0 0 mushroom 22 $\bm{5}$ 0 0 0 0 connect 22 $\bm{5}$ 0 0 0 0 connect 33 $\bm{2}$ 11 11 11 0 munin1 22 $\bm{5}$ 0 0 0 0 andes 22 $\bm{5}$ 0 0 0 0 andes 33 $\bm{5}$ 0 0 0 0
 Table 3: Results of score maximization.
@@
 D.2 Identification of Candidate Parent Sets
-[] Instance $m$ Number Time [s] alarm $4$ $2291\pm 210$ $1700\pm 11$ barley $4$ $355\pm 12$ $235015\pm 95243$ hailfinder $4$ $692\pm 16$ $90247\pm 3334$ hepar2 $4$ $689\pm 59$ $40216\pm 534$ chess $2$ $32570\pm 1686$ $59\pm 0$ chess $3$ $236285\pm 15636$ $5463\pm 284$ win95pts $2$ $9231\pm 665$ $60\pm 3$ win95pts $3$ $43657\pm 3850$ $2404\pm 51$ pathfinder $2$ $50962\pm 3254$ $320\pm 5$ pathfinder $3$ $398131\pm 47889$ $44991\pm 3426$ mushroom $2$ $273526\pm 12647$ $689\pm 26$ connect $2$ $76996\pm 4381$ $331\pm 7$ connect $3$ $710684\pm 73964$ $83361\pm 6624$ munin1 $2$ $135570\pm 6765$ $2466\pm 29$ andes $2$ $7424\pm 219$ $852\pm 52$ andes $3$ $14195\pm 556$ $104683\pm 3855$
+[] Instance $m$ Number Time [s] alarm 44 $2291\pm 210$ $1700\pm 11$ barley 44 $355\pm 12$ $235015\pm 95243$ hailfinder 44 $692\pm 16$ $90247\pm 3334$ hepar2 44 $689\pm 59$ $40216\pm 534$ chess 22 $32570\pm 1686$ $59\pm 0$ chess 33 $236285\pm 15636$ $5463\pm 284$ win95pts 22 $9231\pm 665$ $60\pm 3$ win95pts 33 $43657\pm 3850$ $2404\pm 51$ pathfinder 22 $50962\pm 3254$ $320\pm 5$ pathfinder 33 $398131\pm 47889$ $44991\pm 3426$ mushroom 22 $273526\pm 12647$ $689\pm 26$ connect 22 $76996\pm 4381$ $331\pm 7$ connect 33 $710684\pm 73964$ $83361\pm 6624$ munin1 22 $135570\pm 6765$ $2466\pm 29$ andes 22 $7424\pm 219$ $852\pm 52$ andes 33 $14195\pm 556$ $104683\pm 3855$
 Table 4: Number of candidate parent sets and execution time for their identification.
-The mean and standard deviation of $5$ trials are presented.
+The mean and standard deviation of 55 trials are presented.
 Empty sets were excluded from the candidate parent sets.
@@
 2:  $\Delta\leftarrow$ the incremental return value of $\bar{H}$ when filipping a bit randomly.
-3:     if a random number between 00 and $1$ is less than $\exp(-\Delta/T_{0}^{1-\mathrm{current~{}time}/\mathrm{time~{}limit}})$then
+3:  if a random number between 0 and $1$ is less than $\exp(-\Delta/T_{0}^{1-\mathrm{current~time}/\mathrm{time~limit}})$then
 4:   accept the bit flip.
@@
 D.5 Identification of Candidate Parent Subsets
-[] Instance $m$ Time [s] alarm $4$ $1\pm 0,50\pm 28,10\pm 1$ barley $4$ $0\pm 0,1\pm 0,10\pm 2$ hailfinder $4$ $0\pm 0,0\pm 0,4\pm 1$ hepar2 $4$ $0\pm 0,0\pm 0,2\pm 0$ chess $2$ $6\pm 0,0\pm 0,83\pm 3$ chess $3$ $639\pm 58,3150\pm 167,3480\pm 315$ win95pts $2$ $2\pm 2,0\pm 0,31\pm 1$ win95pts $3$ $49\pm 8,530\pm 73,457\pm 66$ pathfinder $2$ $18\pm 2,0\pm 0,241\pm 29$ pathfinder $3$ $3844\pm 750,2132\pm 129,24731\pm 3458$ mushroom $2$ $277\pm 19,1\pm 0,403\pm 14$ connect $2$ $26\pm 3,0\pm 0,203\pm 15$ connect $3$ $6219\pm 1396,4568\pm 177,19595\pm 2743$ munin1 $2$ $78\pm 6,0\pm 0,875\pm 71$ andes $2$ $1\pm 0,0\pm 0,247\pm 37$ andes $3$ $4\pm 0,7\pm 1,493\pm 131$
+[] Instance $m$ Time [s] alarm 44 $1\pm 0,50\pm 28,10\pm 1$ barley 44 $0\pm 0,1\pm 0,10\pm 2$ hailfinder 44 $0\pm 0,0\pm 0,4\pm 1$ hepar2 44 $0\pm 0,0\pm 0,2\pm 0$ chess 22 $6\pm 0,0\pm 0,83\pm 3$ chess 33 $639\pm 58,3150\pm 167,3480\pm 315$ win95pts 22 $2\pm 2,0\pm 0,31\pm 1$ win95pts 33 $49\pm 8,530\pm 73,457\pm 66$ pathfinder 22 $18\pm 2,0\pm 0,241\pm 29$ pathfinder 33 $3844\pm 750,2132\pm 129,24731\pm 3458$ mushroom 22 $277\pm 19,1\pm 0,403\pm 14$ connect 22 $26\pm 3,0\pm 0,203\pm 15$ connect 33 $6219\pm 1396,4568\pm 177,19595\pm 2743$ munin1 22 $78\pm 6,0\pm 0,875\pm 71$ andes 22 $1\pm 0,0\pm 0,247\pm 37$ andes 33 $4\pm 0,7\pm 1,493\pm 131$
 Table 5: Execution time for IP only. Displays are the same as table 2.